Two cooling fans are installed in some laptop computers. Suppose the reliability of each cooling fan is 0.86. What percent improvement in reliability does adding the second fan provide? Round your answer to the nearest whole number. _________
Answer #1

Given :

Two cooling fans

Reliability = 0.86

As it shows a parallel system, use the formula :

Reliability = 1 - ( 1 - 0.86 ) (1 - 0.86 )

= 1 - ( 0.14 ) ( 0.14 )

= 1 - 0.0196

= 0.9804

Percentage improvement = 1 - 0.9804 / 0.86

= - 0.14

= 14 %

Hence,

The percent improvement in reliability by adding the second fan is 14%